OpenStudy (anonymous):

A spring has a natural length of 20cm. If a 25-N force is required to keep it stretched to a length of 30 cm, how much work is required to stretch it from 20 cm to 25 cm?

OpenStudy (lgbasallote):

the formula is \[F = kx\] 25 N is requred to stretch the spring by 10 cm \[25 = k(10)\] \[k = \frac{25}{10}\] now..how much force is required to stretch it by 5 cm.. \[F = kx\] \[F = (\frac{25}{10}) (5)\] \[F = \frac{25}{2}\] does that help?

OpenStudy (anonymous):

Ok, that gives me the force which I have computed up until then, however, I get caught on when we need to use that to get the total amount of work done on the spring.

OpenStudy (lgbasallote):

oh..you're looking for work if i remember right..spring work is \[W = \frac 12 kx^2\] does that help?

OpenStudy (anonymous):

Yes much thanks. I get a total answer of 0.31. Thank you very much!

OpenStudy (lgbasallote):

welcome